OpenStudy (anonymous):

[100^0(x^-1y^3)^-1/x^2y]^-3

OpenStudy (anonymous):

Use the properties of exponents to simplify each of the following expressions, writing your answer with only positive exponents.

OpenStudy (anonymous):

\[\frac{\left(\frac{y^3}{x}\right)^{\frac{3}{x^2}}}{y^3} \]

OpenStudy (anonymous):

is that the answer? or were you typing the equation?